OpenStudy (icalibear):

is (-2c)^3 simplified -2c^3

OpenStudy (anonymous):

no it's not

OpenStudy (tkhunny):

No. Generally, \((ab)^{n} = a^{n}\cdot b^{n}\) In your particular problem, \((-2c)^{3} = (-2)^{3}\cdot c^{3} = -8c^{3}\)

OpenStudy (icalibear):

okay thanks i just wasnt sure if you cubed the -2

OpenStudy (whpalmer4):

no, exponentiation happens before subtraction or unary minus. you cube the 2, then stick the minus sign on the front.

OpenStudy (anonymous):

if have trouble write it out for example -2c *-2c *-2c

OpenStudy (whpalmer4):

this is a mistake that is made very frequently, even by experienced people who aren't careful!

OpenStudy (icalibear):

thanks guys

OpenStudy (whpalmer4):

sorry, I was a bit unclear: I was referring to the -a^b case: you raise a to the b power, then put the minus sign on the front.

OpenStudy (icalibear):

how would you simplify (v/w^-2)^-2 then?? :s do you flip it orr

OpenStudy (whpalmer4):

I would work from the inside out: v/w^-2 would become vw^2 then (vw^2)^-2 you can just multiply the exponents giving you v^-2w^-4 or 1/(v^2w^4)

OpenStudy (whpalmer4):

or you could go (v^-1w^-2)^2 = v^-2w^-4 = 1/(v^2w^4) just pick a method that works reliably for you and stick with it

OpenStudy (whpalmer4):

shortest path to the answer isn't the best if you make more mistakes!

OpenStudy (anonymous):

but v^-2w^-4 is not simplied at least not where i go to school

OpenStudy (icalibear):

yeah it cant have negative exponents

OpenStudy (anonymous):

yup that's what I'm taught

OpenStudy (icalibear):

i thought you had to flip the fraction because of the negative exponent outside the parenthesis but what about the negative exponent in the denominator????

OpenStudy (tkhunny):

Unary minus is OFTEN of higher precedence than exponentiation. It does depend on the working environment. As a very good general rule, check before you write code. As far as algebra one rules of precedence, I've seen it both ways in different reliable textbooks. Personally, I prefer the unary minus with higher precedence. Obviously, not everyone agrees. It is a mistake to treat unary minus higher ONLY if it is clearly defined contrary to this position. If it is not clearly defined, I am aware of no authoritative guidance one way or the other.

OpenStudy (icalibear):

I dont know then

OpenStudy (anonymous):

you can do these problems a few different ways but the answer at the end is the same. u don't have to do a certain step first... u just need to find what works best for u that results in not making silly errors

OpenStudy (icalibear):

would i flip the fraction so it becomes (w^2/v)^2?

OpenStudy (anonymous):

ur problem isn't (v^-2/w^-2) so if u flipped ur problem inside first it would be (vw^2)^-2

OpenStudy (icalibear):

the outside exponent is actually a -3 ugh sorry

OpenStudy (anonymous):

when doing these types of problems u might want to rewrite it like (v^1w^-2)^-3

OpenStudy (icalibear):

why get rid of the fraction bar?

OpenStudy (icalibear):

sorry i just am having troubles with these

OpenStudy (anonymous):

sorry my bad (v^1/w^-2)^-3

OpenStudy (icalibear):

but to simplify?

OpenStudy (anonymous):

me personally i write mine out like V^(1*-3) / w^(-2*-3)

OpenStudy (icalibear):

and that's simplified?

OpenStudy (anonymous):

no just one step closer ..... then V^-3/w^6 then 1/v^3w^6

OpenStudy (anonymous):

lot times problems get larger so doing it that way helps....

OpenStudy (icalibear):

ok. why the 1? i thought uou jut flipped the negative exponent to the denominator to get w^6 / v^3

OpenStudy (anonymous):

how did u get w^6 on top?

OpenStudy (tkhunny):

\((v/w^-2)^-2 = \dfrac{1}{\left(v/w^{-2}\right)^{2}} = \dfrac{1}{\left(v\cdot w^{2}\right)^{2}} = \dfrac{1}{v^{2}w^{4}}\)

OpenStudy (icalibear):

I thought you always switched the negative exponent to theother side to make it positive

OpenStudy (tkhunny):

\( (v^{1}/w^{-2})^{-3} = \left(v\cdot w^{2}\right)^{-3} = \dfrac{1}{\left(v\cdot w^{2}\right)^{3}} = \dfrac{1}{v^{3}\cdot w^{6}}\)

OpenStudy (anonymous):

if u have 1/w^-2 u can switch that but if u have have v^1/w^-2 u can't switch both only equals v^1w^2

OpenStudy (icalibear):

okay >.<

OpenStudy (tkhunny):

\(\dfrac{v}{w^{-2}} = v\cdot w^{2}\)

OpenStudy (anonymous):

when u have v^-1/ w^-2 u switch both w^2/v^1

OpenStudy (anonymous):

u don't have to write v^1 i just like to show that because it helps seeing it better for some people

OpenStudy (anonymous):

do u think u understand it better now

OpenStudy (icalibear):

kind of